Pre-Lecture Question 05 To answer this question, first click here to view the associated video The 110 V provided by common wall sockets in the United States is given as a root-mean-square (rms) voltage. What is the peak value of the vcltage? Vo = 78.0 V Vo = 110 v Vo 155 v Suppose the rms voltage is 110 V in a single-loop circuit that contains a resistor R-22 Ω, what is the average power dissipated by the resistor in this circuit? 550 W 776 W 2420 W 390 W
Answer #1

Since Vrms = V peak / √2

=) V peak = Vrms ×√2 = 110√2 = 155V

Option third is correct.

Avg power = V^2 /R = 110×110/22 = 550w